The difference of two numbers is 13. If the smaller number is 2 more than one fourth the larger,what are the numbers
Answer by srinimath(2)

The difference of two numbers is 13. If the smaller number is 2 more than one fourth the larger,what are the numbers

Let the numbers be a and b
then we know a - b = 13 and
b = 1/4 (a) + 2
so we can plug one equation in other and solve
a - (1/4 a + 2 ) = 3/4 a - 2 = 13
3/4 a = 15
a = 4 /3 * 15 = 20
so a is 20
b is 7
I am thrilled by registering to www.wisechalk.com, they pay $20 per hour.